In early 2010, one U.S. dollar bought U92.72. In late 2019, it bought about U108.57.

(a) What was the yen price of one U.S. dollar in 2010? And in 2019?

(b) What was the U.S. dollar price of one yen in 2010? And in 2019?
Answer #1

(a) Yen price of one U.S. dollar in 2010 = 92.72 Yen. [Yen/USD]

Yen price of one U.S. dollar in 2019 = 108.57

(b) The U.S. dollar price of one yen in 2010 = 1/92.72 = 0.01078515962 U.S. dollar

The U.S. dollar price of one yen in 2019 = 1/108.57 = 0.009210647509 U.S. dollar

Can you please upvote? Thank You :-)